What is the capacitance of a capacitor that stores 4.2 μC of charge on its plates when a voltage of 1.6 V is applied between them?
Answer #1

Charge q =4.2 x 10-6 C

Voltage V= 1.6 V

using the formula, q = CV

capacitance C =q / V = 4.2 x 10-6 / 1.6 = 2.625 x 10-6 F = 2.625 \muF